OpenStudy (anonymous):

what percent of 67 is 33 ? if necessary, round to the nearest tenth of a percent A.203.0% B.0.5% C.49.3% D. 33.0%

OpenStudy (anonymous):

C.

OpenStudy (ganpat):

x % of 67 == 33 SO, 67 * x/100 = 33 x = 33 * 100 / 67 x = 49.25 % so option C...